A 50kg woman is on large swing of radius 9m that would rotate in vertical circle at 6 rev/min.what is the magnitude of her weight when she has moved halfway up?

Grade:12

1 Answers

Vikas TU
3 years ago
N−mg=−mv^2/r 
To feel weightless at top , N=0
mg=mv^2/r
⇒v=√rg= √90  m/s
…...................................................